The physician orders Digoxin 0.375mg. You have available from Pharmacy 125mcg tablets. How many tablets are...

  1. The physician orders Digoxin 0.375mg. You have available from Pharmacy 125mcg tablets. How many tablets are you going to administer?

  1. The physician orders Amoxicillin 20mg IVP Q6h. Your pharmacy supplies 10mg per 5 ml. How many milliliters will you administer per dose?

  1. Your patient is anxious and you are going to give Ativan to calm her. The physician orders 6 mg IVP and you have 4mg/ml vials. How many ml will you administer?

  1. You evaluate a person’s intake for 12 hours. What is your patient’s intake for your shift?

Intake

             3 cups of coffee

             6 ounces of OJ

             240 ml of soup

             IV fluids 100ml/hr

  1. Using the patient in the last question, what is her output for your 12 hour shift?

Output

Urine- 800m

Chest tube 200 ml for your shift

  1. A patient has a liter bag of fluids to be infused over 12 hours. How many ml per hour should you set your pump to ensure the whole liter is infused?
Answer #1

Ordered dose= 0.375mg

Convert to mcg( multiply by 1000)= 375 mcg

Available dose= 125 mcg

Number of tablet to administer= 375/125=3

Ordered dose of amoxicillin= 20 mg

Available dose= 10 mg per ml

Millilitres to be administered=20/10= 2ml

Ordered dose of ativan=6mg

Available dose=4mg/ml

Medicine to be administered in millilitres=6mg/4mg=1.5ml

1cup cofee=250ml

So, 3 cups=3*250=750ml

6 ounce orange juice=3/4 cup=175=ml

Soup=240 ml

Iv fluids= 100ml per hour

For 12hour shift, 12*100=1200ml

Total fluid intake=750+175+240+1200=2365 ml

Urine output=800ml(urine)+ 200ml(chest tube))=1000ml

1liter bag=1000ml

Time taken to infusion=12 hours

Infusion rate per hour=1000/12=83.33ml per hour